OpenStudy (anonymous):

a silver ball weighing 50g is removed from a furnace and dropped into a 350g of water at 0 degree celcius. if the equilibrium temperature is 10 degree celcius and the specific heat capacity of platinum is 230Jkg-1 degree celcius-1, what is the temperature of the furnace? neglect the effect of the mass of the calorimeter...

OpenStudy (anonymous):

anyone can help me??

OpenStudy (anonymous):

\[Q=mCv DeltaT\] If Qw=Qs

OpenStudy (anonymous):

\[Ms Cs (10-Tf)=Mw Cw 10\]

OpenStudy (anonymous):

Ms mass silver Mw mass water Cs and Cw specific heat capacity of the silver and Water Tf Temperature of the furnace

OpenStudy (anonymous):

10 means what?? ten??

OpenStudy (anonymous):

sorry is 283 K

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

283 K is the equillibrium temperature

OpenStudy (anonymous):

indeed it is ten

OpenStudy (anonymous):

(0.05)(230)(10-Tf)=(0.35)(4200)(10) ??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i get -1268.26

OpenStudy (shane_b):

The question first says the ball is made of silver then it later gives the specific heat of platinum to be 230J/kgC...which is actually the specific heat of silver. At any rate, I'll use the value the question provides for the ball. The only formula you need for this problem is: \[Q=mC \Delta T\] First, calculate the energy required to raise the temperature of the water by 10 degrees C: \[Q_{water}=(350g)(4.186J/gC)(10)=14.651kJ\] For this problem, we can assume all of this energy came from the ball: \[Q_{ball}=Q_{water}=14.651kJ\] Now calculate the change in temperature of the ball: \[Q_{ball}=(50g)(0.23J/gC)(\Delta T)=14.651kJ\]\[\Delta T = 1274C=T_{initial}-T_{final}\]And we know Tfinal is 10C, so that becomes: \[1274C+10C=T_{initial}=1284C\]Therefore, the temperature of the furnace must have been 1284C.

OpenStudy (anonymous):

thanks you...now i understand this question...;)

OpenStudy (shane_b):

no problem :)

OpenStudy (anonymous):

thanks a lot...